# C: A Coverage Path Planning Algorithm for Unknown Environments using Rapidly Covering Graphs

### v1

@@ V-C2 Coverage using a Multi-Robot Team
The performance of C is compared with baseline algorithms (Dec-PPCPP [45] and Artificial Potential Field Coverage Path Planning (APF-CPP) algorithm [46]). Dec-PPCPP is extended from PPCPP [13] into a decentralized multi-robot CPP by adding dynamic predator (other robots) avoidance reward in the reward function. Wang et al. [46] proposed the APF-CPP algorithm which leverages the concept of artificial potential field for decision making of each robot. Fig. 17(a) shows the trajectories of C, APF-CPP and Dec-PPCPP. Fig 17(b) shows the comparative evaluation results in terms of the different performance metrics. It is seen that C provides clean coverage pattern, with minimum coverage time and overlappings, and leaving behind no coverage holes. Overall, C outperforms the baseline algorithms in terms of all metrics.
@@ VI Conclusion and Future Work
The paper presents a novel algorithm, called C, for adaptive coverage path planning in unknown environments. C is built upon the concepts of progressive sampling, incremental RCG construction and TSP-based coverage hole prevention. It is shown that C is computationally efficient and guarantees complete coverage. The comparative evaluation with existing algorithms shows that C significantly improves the coverage time, number turns, trajectory length, and overlap rate. Finally, C is validated by real experiments and also extended to energy-constrained and multi-robot coverage applications. Future research areas include problems related to adaptive CPP for 1) resilient multi-agent coverage considering failures [56], 2) time-risk optimization [57] considering dynamic obstacles, and 3) spatio-temporally varying environment currents [58].

### v2

@@ V-C2 Coverage using a Multi-Robot Team
The performance of C is compared with baseline algorithms (Dec-PPCPP [hassan2020dec] and Artificial Potential Field CPP (APF-CPP) algorithm [wang2024apf]). Dec-PPCPP is extended from PPCPP [hassan2019ppcpp] into a decentralized multi-robot CPP by adding dynamic predator (other robots) avoidance reward in the reward function. Wang et al. [wang2024apf] proposed the APF-CPP algorithm which leverages the concept of artificial potential field for decision making of each robot. Fig. 17(a) shows the trajectories of C, APF-CPP and Dec-PPCPP. Fig 17(b) shows the comparative evaluation results in terms of the different performance metrics. It is seen that C provides clean coverage pattern, with minimum coverage time and overlappings, and leaving behind no coverage holes. Overall, C outperforms the baseline algorithms in terms of all metrics.
@@ VI Conclusion and Future Work
The paper presents a novel algorithm, called C, for adaptive CPP in unknown environments. C is built upon the concepts of progressive sampling, incremental RCG construction and TSP-based coverage hole prevention. It is shown that C is computationally efficient and guarantees complete coverage. The comparative evaluation with existing algorithms shows that C significantly improves the coverage time, number of turns, trajectory length, and overlap rate. Finally, C is validated by real experiments and also extended to energy-constrained and multi-robot coverage applications.
Future research areas include advanced CPP problems that consider 1) decentralized multi-agent coverage for robustness to failures [Hare_POSER2021] and scalability and 2) time-risk optimization [Songgupta2019] of coverage trajectory for environments with dynamic obstacles and spatio-temporally varying currents [mittal2020rapid].
-A Additional Results
This section presents the detailed results for scenarios 3 (forest), 4 (office), 5 (warehouse), and 6 (maze) of Fig. 12(a). Fig. 19 shows the comparative evaluation results of the above scenarios, including the trajectory plots and performance metrics. The results demonstrate that C consistently delivers superior performance in terms of coverage time, number of turns, trajectory length, and overlap rate in all these scenarios.
Figure 19: Performance comparison of C with the baseline algorithms.
-B Scalability Analysis
This section presents the scalability analysis with respect to the map size. If the map size is increased while preserving its structure (i.e., obstacle distribution), then C generates the same trajectory pattern with the same overlap rate. Thus, the trajectory length increases in proportion to the map size. As such, the coverage time increases linearly. To validate this, we conducted a simulation study of varying map sizes. Specifically, we generated three maps with identical obstacle layouts but different dimensions: $25\,\mathrm{m}\times 25\,\mathrm{m}$, $50\,\mathrm{m}\times 50\,\mathrm{m}$, and $100\,\mathrm{m}\times 100\,\mathrm{m}$, as shown in Fig. 20(a). The C trajectories for these three maps are shown in Fig. 20(a), while the performance metrics are shown in Fig. 20(b). As expected, the coverage time and trajectory length increase linearly with the map size.
Figure 20: Scalability analysis of C for different map sizes: (a) coverage paths in maps of $25\times 25\,\mathrm{m}$, $50\times 50\,\mathrm{m}$, and $100\times 100\,\mathrm{m}$ with identical obstacle layouts; (b) performance comparison in terms of coverage time, number of turns, trajectory length, and overlap rate.